# Periodic Asynchrony: An Effective Method for Accelerating Reinforcement Learning for Large Language Models

### v3

@@ Abstract
Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention, with growing efforts to reproduce and apply it. However, training efficiency remains a critical challenge. In mainstream RL frameworks, inference and training are typically deployed on the same devices. While this approach reduces costs through resource consolidation, its synchronous execution imposes a computational coupling that prevents concurrent inference and training. In this study, we are returning to the strategy of separating inference and training deployment, and by introducing improvements in the data loader, we transform the conventional synchronous architecture into a periodically asynchronous framework, which allows for demand-driven, independent, and elastic scaling of each component, while the accuracy of the algorithm remains completely equivalent to the synchronization method, with both belonging to the on-policy strategy. It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation. In practice, these works have achieved at least a threefold overall performance improvement in RL training on NPU platforms, indicating its potential for widespread application.
@@ 1 Introduction
Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations. Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training. Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed. This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm. The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation. Experiments show that on a 16-NPU setup training an 8B model with 16k context, our approach delivers over a threefold improvement in end-to-end RL throughput.
@@ 2 Related Work
Most recently, AReaL Fu et al. (2025) advanced the application of asynchronous RL to LLM reasoning tasks by fully decoupling generation from training. It introduces a parameter $\eta$ to control data staleness, enabling better handling of delayed samples. To better handle delayed samples, it incorporates a staleness-enhanced PPO variant, effectively modifying the on-policy paradigm to tolerate partially off-policy data. Additionally, its training start time depends on the completion of the first inference batch. In a similar vein, ROLL Flash Lu et al. (2025), also recently introduced, shares conceptual similarities with AReaL. Like AReaL, it decouples training from inference and controls the proportion of off-policy data through an async ratio. Both AReaL and ROLL Flash are off-policy asynchronous frameworks that improve training efficiency but require modifications to standard on-policy RL algorithms to handle data bias. This design trades a small portion of data originally intended for strictly on-policy training to avoid inference-service blocking, effectively shifting it toward off-policy learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
In contrast, our method takes a different perspective with several innovations:
@@ 4.3 Periodic Asynchronous Reinforcement Learning
16:     Compute loss $L(s_{i})$ then backward
@@
21:  Update policy’s weights using gradient $O$
@@
Therefore, in theory, when a training batch cannot be processed within a single inference batch, the asynchronous system can achieve more than twice the startup speed of the synchronous system.
@@ 5 Implementation
Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture. We incorporate parts of Megatron-Core v0.8.0 and DeepSpeed v0.17.2 Rasley et al. (2020). On the NPU platform, we additionally use MindSpeed v2.1.0_core_r0.8.0, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023).
For inference, we use vLLM, with ascend_vLLM 0.9.2.rc1 on NPUs. As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
@@ 6 Experiments
Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.725 0 0 0
MindSpeed-RL 100 0.733 706.355 198.942 16 61.641
Sync (ours) 100 0.746 827.306 143.678 16 99.966
Async (ours) 100 0.758 917.609 82.861 16 192.259
Table 1: Accuracy and training efficiency comparison on Qwen3-8B. All frameworks were trained for 100 steps, and training time and token counts were extracted from logs. For our synchronous and asynchronous methods under the training–inference decoupled architecture, 8 NPUs were used for one training instance ($\text{TP}=8,\text{PP}=1$) and 8 NPUs for four inference instances ($\text{TP}=2$). MindSpeed-RL used all 16 NPUs for two training instances ($\text{TP}=8,\text{PP}=1$) and eight inference instances ($\text{TP}=2$).
Framework Training Steps AIME24 Trained Tokens (M) Training Hours NPUs Tokens/Sec/Device
Baseline 0 0.726 0 0 0
MindSpeed-RL 20 0.696 107.874 70.647 64 6.627
Async (ours) 20 0.737 123.999 21.453 48 33.449
Table 2: Accuracy and training efficiency comparison on DeepSeek-R1-Distill-Qwen-32B. All frameworks were trained for 20 steps, and training time and token counts were taken from logs. Our asynchronous setup uses 32 NPUs for one training instance ($\text{TP}=8,\text{PP}=4$) and 16 NPUs for four inference instances ($\text{TP}=4$). MindSpeed-RL uses all 64 NPUs for two training instances ($\text{TP}=8,\text{PP}=4$) and 16 inference instances ($\text{TP}=4$).
This study conducts comparative experiments on mathematical reasoning tasks, evaluating three reinforcement learning frameworks: MindSpeed-RL, which uses a shared-resource architecture for training and inference; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework. The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system. The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used. Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.
@@ 6.1 Model and Environment Configuration
The models compared in this section are Qwen3-8B  Team (2025) and DeepSeek-R1-Distill-Qwen-32B  DeepSeek-AI (2025) (since the current version of Mindspeed-RL does not yet support reinforcement learning training for Qwen3-32B, we instead use DeepSeek-R1-Distill-Qwen-32B as the training model.). The training data comes from the publicly available math problem dataset DeepScaler Luo and others (2025). All reinforcement learning experiments use the GRPO algorithm. The accuracy test set is the AIME24 Zhang and Math-AI (2024).
6.2 Comparison between Asynchronous and Synchronous Training
@@
Given that the prompts for the samples are relatively short while the responses are longer, Shared-Prompt Attention is disabled in the experiments. In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. As shown in Table 1, in terms of TPSPD, the training speed ratio among the three implementations was $\mathbf{1\!:\!1.62\!:\!3.12}$. The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B, with a context length of 16k, and only two experiments were conducted. As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments. According to the results, for the 32B-scale model, the ratio of the TPSPD metric between Mindspeed-RL and our asynchronous framework is $\mathbf{1\!:\!5.05}$.
@@
For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect. The tests were conducted using the inference en gine vLLM 0.9.1.rc with an output length limit of 32k. Each test sample was evaluated 8 times, with inference parameters set to temperature = 0.7, top_p = 0.95, and top_k = 20. The results show that the models trained by all three frameworks achieved similar accuracy on the AIME24 benchmark, with any differences falling within the expected range of random fluctuations. This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
@@
We also monitored the reward scores during training for all three frameworks, as shown in Figure 5(a). The reward trajectories of our synchronous and asynchronous methods are almost identical, suggesting that the asynchronous approach is computationally equivalent to the synchronous one. Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data. Figure 5(b) presents a comparison of reward scores between Mindspeed-RL and our asynchronous architecture, and it appears that the trends are generally consistent.
(a) Synchronous vs. Asynchronous.
(b) Mindspeed-RL vs. our Asynchronous.
@@
Combining the results of the two experimental groups, we obtain a clear comparative plot of the training speeds, as shown in Figure 6.
Figure 6: TPSPD: MindSpeed-RL vs. Async (ours).

### v4

# Periodic Asynchrony: An On-Policy Approach for Accelerating LLM Reinforcement Learning

@@ Abstract
Since the introduction of the GRPO algorithm, reinforcement learning (RL) has attracted increasing attention, with growing efforts to reproduce and apply it. However, training efficiency remains a critical challenge. In mainstream RL frameworks, inference and training are typically deployed on the same devices. While this approach reduces costs through resource consolidation, its synchronous execution imposes a computational coupling that prevents concurrent inference and training. In this study, we are returning to the strategy of separating inference and training deployment, and by introducing improvements in the data loader, we transform the conventional synchronous architecture into a periodically asynchronous framework, which allows for demand-driven, independent, and elastic scaling of each component, while the accuracy of the algorithm remains completely equivalent to the synchronization method, with both belonging to the on-policy strategy. It is worth emphasizing that we apply a unified tri-model architecture in the training phase, and we also proposed a shared-prompt attention mask to reduce repetitive computation. In practice, our approach consistently delivers significant end-to-end training efficiency improvements on NPU platforms, indicating its potential for widespread application.
@@ 1 Introduction
Motivated by these limitations, we revisit the training–inference separation strategy and introduce a synchronous framework enhanced with new system-level optimizations. Our design uses a Megatron-style 3D parallel backbone and adds a background producer process that continuously dispatches prompts to inference workers, while the main consumer process retrieves outputs and performs training. Samples from each group are split into micro-batches and trained sequentially, and model weights are synchronized only after the full batch is consumed. This periodic asynchrony achieves purely on-policy asynchronous acceleration and is compatible with any on-policy RL algorithm. The decoupled architecture enables flexible scaling of inference workers, avoiding inference bottlenecks, and for GRPO, we further introduce a shared-prompt attention mask to eliminate redundant computation. Experiments show that when training a 7B-scale model on 16 NPUs, our proposed optimizations achieve over a twofold improvement in end-to-end RL training throughput.
@@ 2 Related Work
Most recently, AReaL Fu et al. (2025) advances asynchronous RL for LLM reasoning by fully decoupling generation from training and introducing explicit staleness control through a parameter $\eta$. To tolerate delayed samples, it adopts a staleness-aware PPO variant that relaxes the strict on-policy assumption. While effective in improving training efficiency, this design modifies standard on-policy objectives and introduces controlled bias, whose generalization across algorithms and data regimes remains theoretically unclear. In parallel, several recent works focus on system-level asynchrony to improve the scalability of LLM post-training. ROLL Flash Lu et al. (2025), Laminar Sheng et al. (2025), and LlamaRL Wu et al. (2025) propose fully decoupled asynchronous execution frameworks that eliminate global synchronization barriers and maximize throughput under variable rollout latency. Trajectory Balance with Asynchrony Bartoldson et al. (2025) further explores decoupling exploration and learning to accelerate post-training. These systems primarily address efficiency from a systems perspective and typically rely on standard PPO-style updates, without explicitly constraining policy staleness or preserving strict on-policy correctness.
For fully asynchronous designs that incorporate off-policy updates, this design effectively trades a small portion of data for off-policy training to prevent transient blocking between inference and learning. Although their experiments demonstrate the effectiveness of this biased setup, whether it generalizes across all RL algorithms and data regimes remains theoretically unproven.
In contrast, our method takes a complementary algorithmic perspective with several innovations:
@@ 4.3 Periodic Asynchronous Reinforcement Learning
16:     Compute loss $L(s_{i})$ and gradient $\nabla L(s_{i})$
@@
21:  Update policy’s weights with accumulated gradient $O$
@@
Therefore, in theory, the asynchronous system can achieve nearly twice the end-to-end training speed of the synchronous system.
@@ 5 Implementation
Our implementation is built on the PyTorch training framework, using a 3D-parallel distributed architecture. We incorporate parts of Megatron-Core and DeepSpeed Rasley et al. (2020). On the NPU platform, we additionally use MindSpeed, torch_npu, and optimized operators such as npu_fusion_attention, an accelerated attention kernel supporting custom masks and serving as a counterpart to flash_attention Dao (2023). For inference, we use vLLM, with ascend_vLLM on NPUs. As in fine-tuning, all RL samples are computed at their native sequence lengths for both forward and backward passes without padding, following dynamic-length training.
@@ 6 Experiments
This study conducts comparative experiments on mathematical reasoning tasks, evaluating four reinforcement learning frameworks: MindSpeed-RL, the official NPU training framework, which uses a Megatron backend and adopts a shared-accelerator design for training and inference; VERL, a currently mainstream reinforcement learning framework, which uses an FSDP backend and supports asynchronous execution; the synchronous reinforcement learning framework implemented in this work under a decoupled training–inference design; and the proposed asynchronous reinforcement learning framework. The main goal is to experimentally compare periodic asynchronous and synchronous training, assess the performance of the proposed system against mainstream synchronous frameworks, and evaluate the scalability of the periodic asynchronous system. The primary evaluation metric is end-to-end training throughput, measured by tokens trained per second per device (TPSPD), calculated as the total number of tokens processed divided by the total elapsed time and the number of devices used. Accuracy metrics are provided for reference to verify algorithm correctness, as no modifications were made to the reinforcement learning algorithm.
@@ 6.1 Model and Environment Configuration
The models compared in this section are Qwen2.5-7B-Instruct Yang et al. (2024), Qwen3-8B Team (2025), and DeepSeek-R1-Distill-Qwen-32B DeepSeek-AI (2025), as the current version of MindSpeed-RL does not yet support reinforcement learning training for Qwen3-32B. The training data comes from the publicly available math problem datasets DeepScaler Luo and others (2025) and GSM8K Cobbe et al. (2021). All reinforcement learning experiments use GRPO algorithm. The accuracy test set is the AIME24 Zhang and Math-AI (2024).
6.2 Asynchronous Training Speed Comparison
@@
Since the prompts in DeepScaler are significantly shorter than the responses, Shared-Prompt Attention is disabled; nevertheless, our method still achieves substantial performance improvements. In the first set of experiments, Qwen3-8B was used as the training model, with a context length of 16k. As shown in Table 1, in terms of TPSPD, the training speed ratio among the four implementations was $\mathbf{1\!:\!2.52\!:\!1.62\!:\!3.12}$. The model used in the second set of experiments is DeepSeek-R1-Distill-Qwen-32B. As shown in Table 2, the evaluation methodology is identical to that used in the first set of experiments. According to the experimental results, for the 32B-scale model, the TPSPD ratio between MindSpeed-RL and our asynchronous framework reaches $\mathbf{1\!:\!5.05}$ under a 16K context length. Due to consistent out-of-memory issues with VERL at 16K, we additionally evaluate VERL under an 8K context length, where the TPSPD ratio between VERL and our asynchronous framework is $\mathbf{1\!:\!1.76}$. This experiment also shows that insufficient context-length configuration can lead to a degradation in model accuracy. We further observe that, under this regime, inference length is a primary factor affecting end-to-end throughput.
Given the increased problem difficulty, DeepScaler represents a regime in which inference is more time-consuming than training. To study the complementary setting where training dominates inference, we additionally conduct experiments on GSM8K using Qwen2.7-7B-Instruct, where both prompts and responses are relatively short. We compare MindSpeed-RL, VERL, and our asynchronous framework under configurations with Shared-Prompt Attention enabled and disabled, and evaluate their end-to-end throughput and accuracy. As shown in Table 3, Shared-Prompt Attention significantly reduces the number of training tokens and leads to improved training throughput, with the corresponding TPSPD ratios of $\mathbf{1\!:0.91\!:\!0.27\!:\!2.19}$.
Setting AIME24 Micro-BS Training Tokens TPSPD
Base model (8B) 0.725
MindSpeed-RL 0.733 1 706.355 M 61.641
VERL (async) 0.746 16 812.084 M 155.521
Sync (ours) 0.746 1 827.306 M 99.966
Async (ours) 0.758 1 917.609 M 192.259
Table 1: All frameworks are trained for 100 steps using 16 NPUs with a global batch size of 32. For GRPO, each group performs 32 rollouts. Since VERL adopts an FSDP backend, the micro-batch size is set to at least 16 and its asynchronous execution is enabled.
Setting AIME24 NPUs GBS MBS Training Tokens TPSPD
Base model (32B) 0.726
MindSpeed-RL 0.717 64 32 1 107.874 M 6.627
Async (ours) 0.737 48 32 1 123.999 M 33.449
VERL (async) 0.617 64 64 64 157.765 M 44.016
Async (ours) 0.675 64 64 1 188.081 M 77.342
Table 2: All frameworks are trained for 20 steps. For GRPO, each group performs 32 rollouts. In the additional group of experiments, the global batch size (GBS) is increased to 64, since VERL requires a minimum micro-batch size (MBS) of 64 when running on 64 NPUs.
Setting GSM8K Micro-BS Training Tokens TPSPD
Base model (7B) 0.801
MindSpeed-RL 0.890 16 72.534 M 199.142
VERL (async) 0.869 16 90.042 M 182.560
Async (ours) 0.912 1 84.192 M 54.463
Async (ours) 0.919 16 60.926 M 435.596
Table 3: All frameworks are trained on 16 NPUs and complete training on the full GSM8K training set at step 116. For GRPO, each group performs 32 rollouts, and the context length is fixed to 1K for all settings. In our asynchronous framework, a micro-batch size of 1 indicates that Shared-Prompt Attention is disabled, whereas a micro-batch size of 16 means that 16 rollouts share a single prompt. Evaluation is conducted on the GSM8K test set.
@@
For accuracy evaluation, we employed a rule-based method where the predicted answer from the policy model is considered correct if it can be accurately extracted and matches the ground-truth answer in the dataset; otherwise, it is deemed incorrect. The tests were conducted using the inference en gine vLLM 0.9.1.rc. The results show that the models trained by all four frameworks achieved similar accuracy, with any differences falling within the expected range of random fluctuations. This supports the equivalence in performance between the asynchronous and synchronous methods implemented in this study.
@@
We also monitored the reward scores during training for all four frameworks, as shown in Figure 5. The reward trajectories of our synchronous and asynchronous methods almost completely overlap, suggesting that the asynchronous approach is computationally equivalent to the synchronous one. Both curves exhibit a slight upward trend, likely because the baseline model was pre-trained on similar data.